a worker pushing a wheelbarrow of mass 50kg with a constant velocity along a horizontal surface whose coefficient of friction is 0.5..what is the force P exerted by the worker if angle is 30 degrees?
1
Expert's answer
2012-08-28T10:14:38-0400
Ff = F * k = (m * g - P * sinA) * k
Ff = P * cosA
m * g * k - P * sinA * k
= P * cosA
m * g * k = P * (cosA + sinA * k)
P = m * g * k / (cosA + sinA
* k)
P = 50 * 9.81 * 0.5 / (0.8660 + 0.5 * 0.5) = 245.25 / 1.116 = 220 N